OpenStudy (anonymous):

which expression defines the given series for seven terms? -13 + (-16) + (-19) +... help me pleaseee:)

OpenStudy (anonymous):

T=-3n-10

OpenStudy (anonymous):

thank youuuu!!:)

OpenStudy (amistre64):

A{n+1} = A{n} - 3 ; A{0}= -13 defines the sequences; and I spose the sum of the functinal equation would define the series right?

OpenStudy (anonymous):

He's more likely to be right than me take his answer instead It depends what expression you normally use

OpenStudy (amistre64):

is should prolly rewrite my recursion as: \[A_{n}=A_{n-1}-3;A_0=-13\]

OpenStudy (amistre64):

\[A_n = -13 -3n\] perhaps?

OpenStudy (amistre64):

yeah, that start at 1 or 0 can be rather indeterminate :)

OpenStudy (anonymous):

I learnt it by starting at \[t _{1}\] with it being\[ t_{n}\]

OpenStudy (anonymous):

But I remember the \[ A_{n+1}\] being another way of writing it